OpenStudy (anonymous):

The very rigorous medical program at a local university has a 30% drop out rate for each year. If the school admits 1,000 freshmen, how many diplomas will need to be ordered 4 years later? (show your work please :) ) 656 diplomas must be ordered 240 diplomas must be ordered 410 diplomas must be ordered 316 diplomas must be ordered The very rigorous medical program at a local university has a 30% drop out rate for each year. If the school admits 1,000 freshmen, how many diplomas will need to be ordered 4 years later? (show your work please :) ) 656 diplomas must be ordered 240 diplomas must be ordered 410 diplomas must be ordered 316 diplomas must be ordered @Mathematics

OpenStudy (anonymous):

I got 240?

OpenStudy (anonymous):

240

OpenStudy (anonymous):

is correct

OpenStudy (anonymous):

what is the passing percentage ?

OpenStudy (stormfire1):

I got 240 also

OpenStudy (stormfire1):

err 240.1

OpenStudy (anonymous):

Did you guys use the formula A = P(1 - r)^t or you did it another way?

OpenStudy (stormfire1):

I simply did (((1000 * .7)*.7)*.7)*.7

OpenStudy (stormfire1):

Since it was only 4 years that seemed the quickest way

OpenStudy (anonymous):

Oh so you cant do it with 1000*.3 *.3 *.3*.3 right?

OpenStudy (anonymous):

pretty much

OpenStudy (anonymous):

\[ 1000 (1 - 0.3)^4 \]

OpenStudy (anonymous):

Oh ok, thank you :)

OpenStudy (stormfire1):

Yea, the formula works just as well

OpenStudy (anonymous):

How about this problem: The number of milligrams of a drug that remains in a patient’s system after t hours is given by the function A(t) = Iert. Juan was given 500 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 6 hours? 133.90 mg 150.60 mg 99.20 mg 111.58 mg I got b

OpenStudy (anonymous):

what is A(t) = Iert ?

OpenStudy (anonymous):

exatly

OpenStudy (anonymous):

Oh its A(t) = l e ^rt

OpenStudy (anonymous):

The e is the e^x on the calculator usually.

OpenStudy (anonymous):

not sure but i think it's d

OpenStudy (anonymous):

r = rate t = time e = 2.71828... I = 500

OpenStudy (anonymous):

again not sure

OpenStudy (anonymous):

d

OpenStudy (anonymous):

How did you get that answer?

OpenStudy (anonymous):

oh wait at a rate of 20% mistake

OpenStudy (anonymous):

Yeah at the rate of 20%

OpenStudy (stormfire1):

Ooops..I put in 5 hours and not 6. Still, I'm getting 1.66........

OpenStudy (anonymous):

egggsactly

OpenStudy (anonymous):

stilll think d

OpenStudy (anonymous):

Yeah, i forgot to mention that also the exponent has to be negative, sorry

OpenStudy (stormfire1):

Lol, that fixes it :) 150.6

OpenStudy (stormfire1):

\[A = 500*e^{-rt}=500mg*e^{-.2*6hr}=150.597106mg = 150.6mg\]

OpenStudy (anonymous):

haha, thank you :) sorry for not saying that part :( but thank you